OpenStudy (anonymous):

whats 16 2/4 - 4/5? Is it 15 and 8/20?

OpenStudy (anonymous):

16 2/4 = 16 10/20 4/5 = 16/20 16 10/20 - 16/20 = 15 14/20

OpenStudy (anonymous):

How do you get the 16 and 10

OpenStudy (anonymous):

1/5 = 4/20 right? then 4/5 = 16/20 1/4 = 5/20 right? then 2/4 = 10/20

OpenStudy (anonymous):

If i leave out the 16 from the 16 2/4 it looks like this \[{2 \over 4} - {4 \over 5} = { {5 \times 2} \over {4 \times 5}} - { { 4 \times 4 } \over {4 \times 5} } = { 10 \over 20 } - { 16 \over 20 }\]

OpenStudy (anonymous):

ohh i get it thanks soooo much!!